OpenStudy (anonymous):

Medal + Fan help please! Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°

OpenStudy (anonymous):

well what is the angle that point makes with the x axis: use arctan(y/x)

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

well if you do that you find its -45o then you can use the distance formula to find the distance the point is from the origin

OpenStudy (anonymous):

right i got -45 what's the distance formula?

OpenStudy (anonymous):

is it these? x = radius×cos(theta) y = radius×sin(theta)

OpenStudy (anonymous):

for this it will be d = sqrt(x^2+y^2)

OpenStudy (anonymous):

ok so that equals 0 @Brandon77

OpenStudy (anonymous):

no d = sqrt( (4)^2 + (-4)^2 ) = sqrt ( 16 + 16 ) = sqrt(32)

OpenStudy (anonymous):

oops my bad put it into calc wrong. what next?

OpenStudy (anonymous):

so one coordinate can be (sqrt(32),-45o)

OpenStudy (anonymous):

how would we find the other one?

OpenStudy (anonymous):

if you have a negative radius then it basically mirrors you line. so if the radius is -sqrt(32) then you can add 180o to your angle and get back to the same line

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

no problem. :)

OpenStudy (anonymous):

just to make sure it should be (-sqrt(32),135o)

OpenStudy (anonymous):

45+180=225 @Brandon77 ?

OpenStudy (anonymous):

but its -45 + 180